OpenStudy (anonymous):

At a track meet, 50 people ran the 100-meter dash. 2 people finished in 11 seconds, 5 people finished in 12 seconds, 8 people finished in 13 seconds, 10 people finished in 14 seconds, 21 people finished in 15 seconds, 2 people finished in 16 seconds, and 2 people finished in 17 seconds. What is the probability distribution for the finish times?

OpenStudy (yanasidlinskiy):

A probability distribution includes ALL the possible events and must add up to 1. Here you have 50 people, so 50 will be the denominator for each probability ... P(11 seconds) = 2/50 = 0.04 P(12 seconds) = 5/50 = 0.10 Do this the same for 13, 14, 15, 16 and 17 seconds. When you are all done, the sum of the 7 probabilities should add up to 1.

OpenStudy (anonymous):

P(11 seconds)= 2/50=0.04 P(12 seconds)=5/50=0.10 P(13 seconds)=8/50=0.16 P(14 seconds)=10/50=0.02 P(15 seconds)=21/50=0.42 P(16 seconds)=2/50=0.04 P(17 seconds)=2/50=0.04

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@YanaSidlinskiy help! =]

OpenStudy (yanasidlinskiy):

Yes. Sometehing like that!!!!:)

OpenStudy (anonymous):

im only getting 82 though when i add them up! I Dont understand what Im doing wrong....

OpenStudy (anonymous):

I dont understand...

OpenStudy (anonymous):

You messed up when you calculated P(14 seconds) try that one again and you'll notice it should be .2

OpenStudy (ranga):

Yes, 10/50 = 0.2

OpenStudy (anonymous):

OMG thanks soo much =]

OpenStudy (ranga):

You are welcome.

OpenStudy (yanasidlinskiy):

Sorry....I was kind of late...lost connection....Anyways though..I'm glad someone finished helping you:)